# GeoLLaVA: Efficient Fine-Tuned Vision-Language Models for Temporal Change Detection in Remote Sensing

### v1

@@ 7 Limitation & Future Work
Reliance on GPT-4o mini for Annotations: We relied on GPT-4o mini for captioning and annotations, which provided reliable ground truth. However, incorporating annotations from other models could lead to a more comprehensive comparison and capture a wider range of nuances in temporal changes. Future work could explore integrating multiple models to enhance annotation diversity and improve the robustness of temporal change detection.
@@
Limited Dataset Due to High Labeling Costs: The high cost of labeling with GPT-4o mini limited us to 100,000 image pairs for training and 6,000 pairs for testing. Larger datasets would improve model performance by providing more diverse examples for better fine-tuning and generalization. Future efforts should focus on creating larger datasets to enhance model training and allow for better generalization across diverse conditions.
@@
Single Dataset: Although our dataset is diverse in terms of classes, locations, and image characteristics, it is still a single dataset. Using multiple datasets with different collection schemas would improve the model’s generalization across various environments and tasks. In the future, expanding the dataset to include multiple data sources with varying temporal resolutions will be crucial to improving model generalization.
@@
Manual Evaluation and Crowdsourcing: We were only able to manually evaluate the test data due to the large volume of training data. Crowdsourcing annotations could enhance captioning verification and quality, though it would come with increased costs. Future work could explore the feasibility of crowdsourcing annotations to further validate and improve the quality of the dataset at scale.
@@
Hardware Limitations: Hardware limitations restricted us to a single GPU. Access to larger, more advanced GPUs would improve model performance by allowing faster processing, larger batch sizes, and the ability to train models with more parameters. Future research could benefit from leveraging more powerful hardware to expedite training processes and accommodate more complex model architectures.
@@
Supervised Learning Only: We only employed supervised learning, which made it easier for the model to learn. While this approach helped achieve reliable results, it limited the model’s ability to generalize to unseen data. Future work could explore the integration of unsupervised or semi-supervised learning approaches, which, though more challenging, could lead to more robust and generalized results, particularly in low-resource settings.
@@
Model Distillation: Techniques like distillation via the teacher-student model were not explored in this work. Distillation could significantly reduce the model’s complexity and computational resource requirements while maintaining performance levels. Future research could focus on applying model distillation techniques to streamline the model, making it more efficient and suitable for real-world applications with limited computational resources.
Appendix A Dataset Supplementary Details
Distribution of the classes in the dataset, split between train and test splits in Figure 5.
Figure 5: Distribution of the classes in the training and testing dataset.
List of templates used to create prompts for fine-tuning the models:
  • "Provide a detailed description of the satellite video, where each frame corresponds to a different time but the same location."
  • "Describe the satellite video thoroughly, noting that each frame shows the same location at a different time.", "Give a detailed account of the satellite video, with each frame depicting the same location at distinct points in time."
  • "Offer an elaborate explanation of the satellite video, where every frame captures the same location but at different times."
  • "Elaborate on the changes in the location as seen in the satellite video, where each frame is a snapshot of the same place at different times."
  • "Provide a report describing the satellite video, where each frame shows the same location at different time points.
For each entry, a random template is used as a prompt for the model, while the API response is used as the answer, thus creating annotations in the required format.
Appendix B Results and Evaluation Supplementary Details
Equations for the ROUGE evaluation metrics are listed here in this appendix.
For ROUGE-1:
$\text{Precision}=\frac{\sum\text{Count}_{\text{match unigrams in output}}}{\sum\text{Count}_{\text{unigrams in output}}}$
$\text{Recall}=\frac{\sum\text{Count}_{\text{match unigrams in output}}}{\sum\text{Count}_{\text{unigrams in reference}}}$
[…]
For ROUGE-2:
$\text{Precision}=\frac{\sum\text{Count}_{\text{match bigrams in output}}}{\sum\text{Count}_{\text{bigrams in output}}}$
$\text{Recall}=\frac{\sum\text{Count}_{\text{match bigrams in output}}}{\sum\text{Count}_{\text{bigrams in reference}}}$
[…]
For ROUGE-L:
the equations use the concept of the Longest Common Subsequence (LCS). The precision, recall, and F1-score can be defined similarly, but are specifically based on the LCS length. While explicit formulas vary by implementation, the fundamental concept involves calculating the LCS between the system-generated output and the reference text, then applying similar formulas for precision, recall, and F1-score as with ROUGE-1 and ROUGE-2 Lin (2004).
$\text{Precision}_{\text{LCS}}=\frac{\text{Length of LCS}}{\text{Length of the system-generated output}}$
$\text{Recall}_{\text{LCS}}=\frac{\text{Length of LCS}}{\text{Length of the reference text}}$
[…]
Parameter Value
MAX_LENGTH 400
MODEL LLaVA-NeXT-Video-7B-hf
USE_QLORA True (4-Bit)
batch_size 2
lora_r 64
lora_alpha 128
Training Configuration
max_epochs 1
val_check_interval 0.2
check_val_every_n_epoch 1
gradient_clip_val 1.0
accumulate_grad_batches 1
learning_rate 1e-4
num_nodes 1
warmup_steps 50
Table 3: Full hyper-parameters used for fine-tuning the final model.

### v2

@@ 7 Limitation & Future Work
Reliance on GPT-4o Mini for Annotations: We relied on a single model for annotations, providing reliable ground truth. However, leveraging other models could offer a broader comparison and capture more nuances in temporal changes. Future work could integrate multiple models to improve annotation diversity and robustness.
@@
Limited Dataset Due to Labeling Costs: Labeling with GPT-4o mini limited us to 100,000 image pairs for training and 6,000 for testing. Larger datasets would provide diverse examples, improving fine-tuning and generalization. Future efforts should prioritize creating larger datasets to enhance model training and performance.
@@
Single Dataset: Our dataset, while diverse in classes, locations, and image characteristics, is still singular. Using datasets with varied collection schemas could improve generalization. Future work should include multiple data sources with different temporal resolutions to expand generalization.
@@
Manual Evaluation and Crowdsourcing: Manual evaluation was restricted to test data due to the training data volume. Crowdsourcing could enhance verification and quality but increases costs. Future work could explore crowdsourcing to validate and improve dataset quality at scale.
@@
Hardware Limitations: We used a single GPU, limiting processing speed and batch sizes. Access to more advanced GPUs would enable faster training, larger batches, and more complex models. Future research should leverage better hardware for training efficiency.
@@
Supervised Learning Only: We used supervised learning, which provided reliable results but limited generalization to unseen data. Future work could explore unsupervised or semi-supervised approaches for more robust results, especially in low-resource settings.
@@
Model Distillation: Teacher-student distillation techniques were not explored. Distillation could reduce model complexity and computational needs while maintaining performance levels. Future research could focus on distillation to make models more efficient for real-world applications.